The displacement of the particle from a reference point is given by √x=2t-3 .Find the distance traveled after 2 sec

Grade:10

1 Answers

vaidik sharma
6 years ago
According to the question;-
x=(2t-3)^2
Differentiating the equation w.r.t. time:-
v=dx/dt=d/dt(2t-3)^2= 2(2t-3)                     [since,x^a=ax^a-1]
we again differentiate  this equation w.r.t. time;-
a=dv/dt=d/dt[2(2t-3)]=4m/s^2.
when t=1, v=-2m/s
t=4/3 , v=0m/s
t=2, v=2m/s
 
V=D/T, thus D=VT;
Therefore, D=2X2=4m
the distance travelled at t=2sec. is 4m.